# On the stability and conditioning of a fictitious domain formulation for fluid-structure interaction problems

### v1

@@ Abstract.
We consider a distributed Lagrange multiplier formulation for fluid-structure interaction problems in the spirit of the fictitious domain approach. Our previous studies showed that the formulation is unconditionally stable in time and that its mixed finite element discretization is well-posed. In this paper, we analyze the behavior of the condition number with respect to mesh refinement. Moreover, we observe that our formulation does not need any stabilization term in presence of small cut cells and conditioning is not affected by the interface position.
@@ 1. Introduction
The discretization of these problems is challenging since it requires an accurate representation of the underlying geometry. To this aim, several methods have been developed during the last decades. They can be classified as fitted or unfitted depending on the fact that the computational mesh is aligned or not with the interface. Among fitted approaches, the most popular is the Arbitrary Lagrangian–Eulerian (ALE) [37, 38, 29]. Fluid and solid domains are discretized by two meshes evolving around a shared interface. Despite the advantage of having the kinematic constraints imposed by construction, strong mesh distortions may appear. Moreover, ALE discretizations of fluid-structure interaction problems where fluid and solid have the same or similar densities may suffer from the added mass effect [23], which causes the instability of semi-implicit time advancing schemes. To mitigate this unpleasant phenomenon, transmission conditions should be treated with appropriate techniques [28, 7].
Unfitted approaches may overcome the above critical issues. In this case, fluid and solid are discretized independently, with meshes not fitting at the interface. Thus, the complexity of mesh generation is minimized. This comes at the price of reducing the accuracy of the scheme unless special techniques are used. Within this family of methods, we mention, for instance, the level set method [42, 24], the immersed boundary method [40, 41], the fictitious domain approach [34, 33], the immersed boundary-conformal isogeometric method [44], the shifted boundary method [39, 6], the virtual element method on polygonal pixel-based tessellations [11], the fat boundary method [10] and, finally, the Nitsche–XFEM method [1, 19].
Our fictitious domain approach [12, 16] originated from the immersed boundary method [17]. The fluid dynamics is studied in a fixed domain, which is fictitiously extended to the region occupied by the immersed structure. In order to describe the solid deformation, we consider a reference domain which is mapped, at each time instant, into the actual configuration. We use a distributed Lagrange multiplier to impose, at variational level, the kinematic constraint describing the interaction between fluid and solid. The finite element discretization of the coupling term is challenging since it requires to integrate functions defined on independent and non-matching grids. Such computation can be carried out exactly by employing a composite quadrature rule on the intersection between the fluid mesh and the mapped solid one. An inexact procedure can also be considered by taking into account the presence of an additional source of error. We studied computational and theoretical aspects regarding the coupling term in our recent works [13, 14].
It is quite common for unfitted finite element methods to require additional stabilization/penalization terms at elements cut by the interface so that well-posedness and accuracy are ensured independently of the interface position, as well as ill-conditioning of the system is avoided. This is the case of Cut–FEM [43, 35, 18, 22], which exploits a suitable penalty term in the interface zone to extend coercivity to the whole background mesh. Inf-sup stability of the fictitious domain method presented in [20, 21] is guaranteed if the jump of the Lagrange multiplier is penalized on element faces. Similar arguments are also considered for the Finite Cell Method [26, 27, 31] and the Nitsche–XFEM mentioned above.
@@
In this work, we recall some properties of our distributed Lagrange multiplier formulation and we prove upper bounds for the condition number. Our method is unconditionally stable in time so that no condition at all is required on the choice of the time step, overcoming the critical issues of the ALE approach. Moreover, the inf-sup stability is guaranteed at discrete level without the need for any artificial penalization term dealing with small cut cells. We then prove upper bounds for the condition number, showing that it only depends on the choice of the coupling term, but not on the position of the interface. Our theoretical results are confirmed by extensive numerical investigations. A spectral analysis has been carried out in [2, 8] to study the effect of preconditioning strategies when solving interface problems using our fictitious domain formulation.
@@
The paper is organized as follows. After recalling the functional analysis notation, we derive our fictitious domain formulation starting from the physical background (see Section 3). In Section 4, we recall the time semi-discretization and the related stability property. The analysis of continuous and discrete problems are reported in Section 5, whereas computational aspects regarding the coupling term are recalled in Section 6. Theoretical results concerning the discrete problem with inexact coupling term are presented in Section 7, before proving the upper bounds for the condition number in Section 8. Our numerical investigation is finally presented in Section 9. The derivation of the model and the results of Sections 5, 6 and 7 are essentially known. We summarized them here since we need them for our analysis of Section 8 with the intention of being self-contained.
2. Notation
@@
Let $D\subset\mathbb{R}^{2}$ be an open bounded domain. The space of square-integrable functions on $D$ is denoted by $L^{2}(D)$ and endowed with the scalar product $(\cdot,\cdot)_{D}$, inducing the norm $\|\cdot\|_{0,D}$. We denote by $L^{2}_{0}(D)$ the subspace of zero mean valued functions. Sobolev spaces are denoted by $W^{s,p}(D)$, with $s\in\mathbb{R}$ being the differentiability exponent and $p\in[1,\infty]$ the integrability exponent. When $p=2$, we adopt the notation ${H^{s}(D)=W^{s,2}(D)}$. The space $H^{s}(D)$ is endowed with the norm $\|\cdot\|_{s,D}$. Finally, the space $H^{1}_{0}(D)\subset H^{1}(D)$ contains functions with zero trace on $\partial D$. Vector spaces and vector/tensor valued functions are denoted by boldface letters.
@@ 3. Problem setting
We focus on fluid-structure interaction problems consisting of an elastic body immersed in a Newtonian fluid. The configuration of the system is described by two regions $\Omega^{s}_{t}$ and $\Omega^{f}_{t}$ representing the position of solid and fluid at time $t$, respectively. More precisely, both $\Omega^{s}_{t}$ and $\Omega^{f}_{t}$ can be either two- or three-dimensional domains. We point out that our approach can also deal with codimension one (see e.g. [12, 5]) and codimension two solids (see e.g. [4, 36]). We then introduce a third domain $\Omega=\Omega^{s}_{t}\cup\Omega^{f}_{t}$, acting as a container for the system and independent of time. We assume that the solid interface $\partial\Omega^{s}_{t}$ does not touch the boundary of the container $\partial\Omega$.
@@
(2) $\rho_{f}\bigg{(}\frac{\partial\mathbf{u}^{f}}{\partial t}+\mathbf{u}^{f}\cdot\nabla\mathbf{u}^{f}\bigg{)}=\operatorname{div}\bm{\sigma}^{f}$ in $\quad\text{in }\Omega^{f}_{t}$
$\operatorname{div}\mathbf{u}^{f}=0$ in $\quad\text{in }\Omega^{f}_{t}$
@@
$|\mathbb{F}|=1$ in .$\quad\text{in }\mathcal{B}.$
@@
$\mathbf{u}^{f}(0)=\mathbf{u}_{0}^{f}\ \text{ in }\Omega^{f}_{0},\qquad\mathbf{u}^{s}(0)=\mathbf{u}_{0}^{s}\ \text{ in }\Omega^{s}_{0},\qquad\mathbf{X}(0)=\mathbf{X}_{0}\ \text{ in }\mathcal{B}.$
(4) $\mathbf{u}=\begin{cases}\mathbf{u}^{f}\quad\text{ in }\Omega^{f}_{t}\\ \mathbf{u}^{s}\quad\text{ in }\Omega^{s}_{t},\end{cases}\qquad p=\begin{cases}p^{f}\quad\text{ in }\Omega^{f}_{t}\\ p^{s}\quad\text{ in }\Omega^{s}_{t}.\end{cases}$
@@ 4. Time discretization and saddle point problem
$\frac{\mathbf{X}^{0}-\mathbf{X}^{-1}}{\Delta t}=\mathbf{u}_{0}^{s}\quad\text{ in }\mathcal{B}.$
We recall that the method is unconditionally stable, without any restriction on the choice of time step $\Delta t$, see [12, Prop. 3].
@@
5. Known theoretical results on continuous and discrete problems
@@
(17) $\mathbf{V}_{h}$ =$=\{\mathbf{v}\in\mathbf{H}^{1}(\Omega):\mathbf{v}_{\mid\mathrm{T}}\in[\mathcal{P}_{1}(\mathrm{T})]^{2}\quad\forall\mathrm{T}\in\mathcal{T}_{h/2}^{\Omega}\}$
$Q_{h}$ =$=\{q\in L^{2}_{0}(\Omega)\cap H^{1}(\Omega):q_{\mid\mathrm{T}}\in\mathcal{P}_{1}(\mathrm{T})\quad\forall\mathrm{T}\in\mathcal{T}_{h}^{\Omega}\}$
$\mathbf{S}_{h}$ =$=\{\mathbf{w}\in\mathbf{H}^{1}(\mathcal{B}):\mathbf{w}_{\mid\mathrm{T}}\in[\mathcal{P}_{1}(\mathrm{T})]^{2}\quad\forall\mathrm{T}\in\mathcal{T}_{h}^{\mathcal{B}}\}$
${\bm{\Lambda}}_{h}$ =$=\{\bm{\mu}\in\mathbf{H}^{1}(\mathcal{B}):\bm{\mu}_{\mid\mathrm{T}}\in[\mathcal{P}_{1}(\mathrm{T})]^{2}\quad\forall\mathrm{T}\in\mathcal{T}_{h}^{\mathcal{B}}\}.$
@@
This discrete problem is well-posed thanks to the theory discussed in [16] combined with [45]. The following proposition holds.
@@
$\leq C\big{(}\inf_{\mathbf{v}_{h}\in\mathbf{V}_{h}}\left\lVert\mathbf{u}-\mathbf{v}_{h}\right\rVert_{1,\Omega}+\inf_{q_{h}\in Q_{h}}\left\lVert p-q_{h}\right\rVert_{0,\Omega}+\inf_{\mathbf{Y}_{h}\in\mathbf{S}_{h}}\left\lVert\mathbf{X}-\mathbf{Y}_{h}\right\rVert_{1,\mathcal{B}}+\inf_{\bm{\mu}_{h}\in{\bm{\Lambda}}_{h}}\left\lVert\bm{\lambda}-\bm{\mu}_{h}\right\rVert_{{\bm{\Lambda}}}\big{)}.$
@@ 6. Computational aspects regarding the coupling term
It is clear that, if the coupling term is assembled with the formulas given in (23), then a quadrature error is produced. This topic has been addressed in [14, Sec. 7]. We report here the associated quadrature error estimates.
@@
7. Results on inexact coupling
@@
We recall the main results from [14] about estimates for the quadrature error as well as the analysis of Problem 6. The well-posedness of Problem 6 is a consequence of the results in [14] combined with the theory in [45]. The following proposition holds true.
@@
In the statement of Problem 6, without affecting the generality of the results, we assumed that the three terms on the right hand side are computed exactly. Actually, the error produced by inexact integration can be easily estimated by standard arguments as in [25, Chap. 4, Sect. 4.1]. The interested reader can find these results in [14, Lemma 1].
@@ 8. Estimate of the condition number
Let $C_{\star},\,C^{\star}$ be two positive constants, independent of the mesh sizes, arising from the equivalence between the $L^{2}$ norm of a finite element function and the Euclidean norm of the corresponding vector of degrees of freedom (see [30, Sect. 2.3]). Then, the following bounds hold
@@ 8.2. Condition number in case of inexact coupling
(41) $\mathcal{L}_{h}(\mathcal{U}_{h},\mathcal{V}_{h};\overline{\mathbf{X}})\leq M\|\mathcal{U}_{h}\|_{\mathbb{V}}\|\mathcal{V}_{h}\|_{\mathbb{V}}\qquad\forall\mathcal{U}_{h},\mathcal{V}_{h}\in\mathbb{V}_{h}.\$
@@ 8.3. Main result
(42) $\mathsf{cond}(\mathcal{A}_{\square})\leq Ch^{-4}$ if $\quad\text{if }\mathbf{c}=\mathbf{c}_{0}\text{ or }\mathbf{c}_{h}=\mathbf{c}_{0,h},$
$\mathsf{cond}(\mathcal{A}_{\square})\leq Ch^{-2}$ if $\quad\text{if }\mathbf{c}=\mathbf{c}_{1}\text{ or }\mathbf{c}_{h}=\mathbf{c}_{1,h}.$

### v2

@@ Abstract.
We consider a distributed Lagrange multiplier formulation for fluid-structure interaction problems in the spirit of the fictitious domain approach. This is an unfitted method, which does not require the construction of meshes conforming to the interface. We focus on the stationary problem arising from the time discretization and we analyze the behavior of the condition number with respect to mesh refinement. At the numerical level, the computation of the term coupling the fluid and the solid mesh requires the knowledge of the intersection between fluid and mapped solid elements and it might happen that a portion of the intersected elements is very small. We show that our formulation is stable independently of such intersections and that the conditioning is not affected by the interface position.
@@ 1. Introduction
The discretization of these problems is challenging since it requires an accurate representation of the underlying geometry. To this aim, several methods have been developed during the last decades. They can be classified as fitted or unfitted depending on the fact that the computational mesh is aligned or not with the interface. Among fitted approaches, the most popular is the Arbitrary Lagrangian–Eulerian (ALE) [33, 34, 24]. Fluid and solid domains are discretized by two meshes evolving around a shared interface. Despite the advantage of having the kinematic constraints imposed by construction, strong mesh distortions may appear. Unfitted approaches may overcome this issue. In this case, fluid and solid are discretized independently, with meshes not fitting at the interface. Thus, the complexity of mesh generation is minimized. In some cases, this comes at the price of reducing the accuracy of the scheme unless special techniques are used. Among unfitted approaches for FSI and multiphysics problems, we mention, for instance, the level set method [38, 19], the immersed boundary method [36, 37], the fictitious domain approach [29, 28, 42], the immersogeometric method [35], the shifted boundary method [21], the fat boundary method [39], and, finally, the Nitsche method. Within the family of Nitsche-type methods, we find schemes based on a fictitious domain approach as in [16, 17], the CutFEM [40, 30, 13, 15], the Finite Cell Method [22, 23, 26] and the Nitsche–XFEM [1, 14].
Our approach falls within the class of unfitted methods for FSI. It originated from the finite element immersed boundary method [10] and further evolved into a fictitious domain method based on the use of a Lagrange multiplier [6, 12]. We consider the simulation of fluid-structure interaction problems which couple an incompressible Newtonian fluid with an incompressible visco-elastic solid body. This setting is motivated by the fact that many biological tissues satisfy this assumption, see e.g. [36]. Actually, our technique can be extended also to the case of compressible solids as in [11]. See also [31], where the finite element immersed boundary method is generalized to an incompressible Newtonian fluid interacting with general hyper-elastic structures.
The main idea is to fictitiously extend the fluid domain to the region occupied by the immersed structure so that we solve the Navier–Stokes equations using the Eulerian framework in a fixed domain. For the solid deformation we use the standard Lagrangian framework and solve the governing elasticity equation on a reference domain mapped, at each time instant, into the actual configuration. A distributed Lagrange multiplier is introduced to deal with the kinematic constraint describing the interaction between fluid and solid at variational level. The fluid and the solid equations are then coupled by additional terms depending on the Lagrange multiplier. In particular, in the fluid equation, this term involves functions defined on the two domains, hence we refer to it as coupling term.
At the computational level, the main advantage of our approach is that we can use totally independent meshes which do not evolve in time. However, the computation of the coupling term requires to integrate on the elements of the reference mesh mapped on the background fluid mesh. Such computation requires the knowledge of the intersections between the fluid mesh and the mapped solid one. Hence, very small cells contained in the support of both fluid and mapped solid basis functions may originate. We refer to them as small cut cells. Then, the matrix associated with the coupling term can be evaluated exactly by employing a quadrature rule on each intersected element. An inexact procedure can also be considered, using a single quadrature rule on each mapped element. In this case, it might happen that no quadrature nodes fall inside a small cut cell. We estimated the error arising from the use of inexact quadrature in our recent works [8, 9]. In this paper, we discuss the effect of the presence of small cut cells on the stability and conditioning of the considered formulation. On the other hand, in the case of Nitsche-type methods, the term cut cell refers to a mesh element which is cut by the interface and it might happen that one portion of the cut cell is small. It is known that this fact can produce instabilities in the solution, as well as ill-conditioning, unless stabilization/penalization terms, such as the ghost penalty technique [18], are added to the variational formulation.
@@
In this work, we recall the main features of our distributed Lagrange multiplier formulation and we prove upper bounds for the condition number, depending only on the choice of the coupling term and not on the integration technique or the interface position. On the other hand, the inf-sup stability is guaranteed at the discrete level without the need for any artificial penalization term dealing with small cut cells.
Our theoretical results on the stability and conditioning are confirmed by extensive numerical investigations. A final numerical test involving a time dependent problem shows that the method is unconditionally stable in time, independently of the presence of small cut cells.
@@
The paper is organized as follows. After recalling the functional analysis notation, we derive our fictitious domain formulation starting from the physical background (see Section 2). In Section 3, we recall the time semi-discretization and the related stability properties, whereas computational aspects regarding the coupling term are recalled in Section 4. We prove the upper bounds for the condition number in Section 5. Our numerical investigation is finally presented in Section 6. The derivation of the model and the results of Sections 3 and 4 are essentially known. We summarized them here since we need them for our analysis of Section 5 with the intention of being self-contained.
@@
Notation. Let $D\subset\mathbb{R}^{2}$ be an open bounded domain. The space of square-integrable functions on $D$ is denoted by $L^{2}(D)$ and endowed with the scalar product $(\cdot,\cdot)_{D}$, inducing the norm $\|\cdot\|_{0,D}$. We denote by $L^{2}_{0}(D)$ the subspace of zero mean valued functions. Sobolev spaces are denoted by $W^{s,p}(D)$, with $s\in\mathbb{R}$ being the differentiability exponent and $p\in[1,\infty]$ the integrability exponent. When $p=2$, we adopt the notation $H^{s}(D)=W^{s,2}(D)$}. The space $H^{s}(D)$ is endowed with the norm $\|\cdot\|_{s,D}$. Finally, the space $H^{1}_{0}(D)\subset H^{1}(D)$ contains functions with zero trace on $\partial D$. Vector spaces and vector/tensor valued functions are denoted by boldface letters.
@@ 2. Problem setting
We focus on fluid-structure interaction problems consisting of an elastic body immersed in a Newtonian fluid. The configuration of the system is described by two regions $\Omega^{s}_{t}$ and $\Omega^{f}_{t}$ representing the position of solid and fluid at time $t$, respectively, and sharing a common interface $\partial\Omega^{s}_{t}\cap\partial\Omega^{f}_{t}$. More precisely, both $\Omega^{s}_{t}$ and $\Omega^{f}_{t}$ can be either two- or three-dimensional domains. We point out that our approach can also deal with codimension one (see e.g. [6, 4]) and codimension two solids (see e.g. [3, 32]). We then introduce a third domain $\Omega=\Omega^{s}_{t}\cup\Omega^{f}_{t}$, acting as a container for the system and independent of time. We assume that the solid domain $\Omega^{s}_{t}$ is totally immersed in the fluid one, so that its closure does not touch the boundary of the container $\partial\Omega$. Therefore, the interface $\partial\Omega^{s}_{t}\cap\partial\Omega^{f}_{t}$ coincides with $\partial\Omega^{s}_{t}$.
@@
(2) $\rho_{f}\bigg(\frac{\partial\mathbf{u}^{f}}{\partial t}+\mathbf{u}^{f}\cdot\nabla\mathbf{u}^{f}\bigg)=\operatorname{div}\bm{\sigma}^{f}$ $\quad\text{in }\Omega^{f}_{t}$
$\operatorname{div}\mathbf{u}^{f}=0$ $\quad\text{in }\Omega^{f}_{t}$
@@
$|\mathbb{F}|=1$ in $\quad\text${in $\mathcal{B}$.
@@
$\mathbf{u}^{f}=\mathbf{u}^{s}$ $\text{ on }\partial\Omega^{s}_{t}$
$\bm{\sigma}_{f}\mathbf{n}_{f}=-\big(|\mathbb{F}|^{-1}\mathbb{T}\mathbb{F}^{\top})\mathbf{n}_{s}$ $\text{ on }\partial\Omega^{s}_{t},$
@@
Remark 1.
By exploiting the incompressibility of both the fluid and solid materials, it is straightforward to impose the divergence free condition to the extended velocity. Notice that, in this case, the variable $p$ represents the Lagrange multiplier associated to such constraint, as usual for the Navier–Stokes equations. This might be considered as a limitation of our method. However, it is possible to extend our fictitious domain approach to the case of compressible solids. This is done for instance in [11], where a fictitious pressure field without any physical meaning is introduced in the solid domain, and it is weakly imposed to be zero. Moreover, in our analysis we strongly rely on the assumption that the solid material has a viscous component. The method can be generalized to hyper-elastic structures as done in [31], where the authors take due account of the additional viscous term that is fictitiously added to the extended fluid equation.
@@ 3. Time discretization and saddle point problem
[…]
We recall that the method is unconditionally stable, without any restriction on the choice of time step $\Delta t$, see [6, Prop. 3].
@@
3.1. Known theoretical results on continuous and discrete problems
@@
(17) $\mathbf{V}_{h}$ $=\{\mathbf{v}\in\mathbf{H}^{1}(\Omega):\mathbf{v}_{\mid\mathrm{T}}\in[\mathcal{P}_{1}(\mathrm{T})]^{2}\quad\forall\mathrm{T}\in\mathcal{T}_{h/2}^{\Omega}\}$
$Q_{h}$ $=\{q\in L^{2}_{0}(\Omega)\cap H^{1}(\Omega):q_{\mid\mathrm{T}}\in\mathcal{P}_{1}(\mathrm{T})\quad\forall\mathrm{T}\in\mathcal{T}_{h}^{\Omega}\}$
$\mathbf{S}_{h}$ $=\{\mathbf{w}\in\mathbf{H}^{1}(\mathcal{B}):\mathbf{w}_{\mid\mathrm{T}}\in[\mathcal{P}_{1}(\mathrm{T})]^{2}\quad\forall\mathrm{T}\in\mathcal{T}_{h}^{\mathcal{B}}\}$
${\bm{\Lambda}}_{h}$ $=\{\bm{\mu}\in\mathbf{H}^{1}(\mathcal{B}):\bm{\mu}_{\mid\mathrm{T}}\in[\mathcal{P}_{1}(\mathrm{T})]^{2}\quad\forall\mathrm{T}\in\mathcal{T}_{h}^{\mathcal{B}}\}.$
@@
This discrete problem is well-posed thanks to the theory discussed in [12] combined with [41]. The following proposition holds true independently of how the fluid and the mapped solid meshes intersect each other.
@@
$\hskip-14.22636pt\leq C\big(\inf_{\mathbf{v}_{h}\in\mathbf{V}_{h}}\left\lVert\mathbf{u}-\mathbf{v}_{h}\right\rVert_{1,\Omega}+\inf_{q_{h}\in Q_{h}}\left\lVert p-q_{h}\right\rVert_{0,\Omega}+\inf_{\mathbf{Y}_{h}\in\mathbf{S}_{h}}\left\lVert\mathbf{X}-\mathbf{Y}_{h}\right\rVert_{1,\mathcal{B}}+\inf_{\bm{\mu}_{h}\in{\bm{\Lambda}}_{h}}\left\lVert\bm{\lambda}-\bm{\mu}_{h}\right\rVert_{{\bm{\Lambda}}}\big).$
@@ 4. Computational aspects regarding the coupling term
It is clear that, if the coupling term is assembled with the formulas given in (23), then a quadrature error is produced. This topic has been addressed in [9, Sec. 7]. We report here the associated quadrature error estimates.
@@
4.1. Results on inexact coupling
@@
We recall the main results from [9] about estimates for the quadrature error as well as the analysis of Problem 6. The well-posedness of Problem 6 is a consequence of the results in [9] combined with the theory in [41]. Also in this case, the following proposition states the stability of the problem without any assumption on the intersection between the fluid and the mapped solid meshes.
@@
In the statement of Problem 6, without affecting the generality of the results, we assumed that the three terms on the right hand side are computed exactly. Actually, the error produced by inexact integration can be easily estimated by standard arguments as in [20, Chap. 4, Sect. 4.1]. The interested reader can find these results in [9, Lemma 1].
@@ 5. Estimate of the condition number
Let $C_{\star},\,C^{\star}$ be two positive constants, independent of the mesh sizes, arising from the equivalence between the $L^{2}$ norm of a finite element function and the Euclidean norm of the corresponding vector of degrees of freedom (see [25, Sect. 2.3]). Then, the following bounds hold
@@ 5.2. Condition number in case of inexact coupling
(41) $\mathcal{L}_{h}(\mathcal{U}_{h},\mathcal{V}_{h};\overline{\mathbf{X}})\leq M\|\mathcal{U}_{h}\|_{\mathbb{V}}\|\mathcal{V}_{h}\|_{\mathbb{V}}\qquad\forall\mathcal{U}_{h},\mathcal{V}_{h}\in\mathbb{V}_{h}.\\$
@@ 5.3. Main result
(42) $\mathsf{cond}(\mathcal{A}_{\square})\leq Ch^{-4}$ $\quad\text{if }\mathbf{c}=\mathbf{c}_{0}\text{ or }\mathbf{c}_{h}=\mathbf{c}_{0,h},$
$\mathsf{cond}(\mathcal{A}_{\square})\leq Ch^{-2}$ $\quad\text{if }\mathbf{c}=\mathbf{c}_{1}\text{ or }\mathbf{c}_{h}=\mathbf{c}_{1,h}.$